23 men shake hand with each other in a party. Maximum no of handshakes without cyclic handshakes
Answer by stanbon(75887)

23 men shake hand with each other in a party. Maximum no of handshakes without cyclic handshakes
-----------------
# of handshakes = 23C2 = (23*22)/(1*2) = 23*11 = 253
===================================
Cheers,
Stan H.